The area of the floor in James' rectangular storage rental locker is 144 sqaure feet. The Length of the locker is 30 feet less than six times the width. What are the dimensions of the floor?
Answer by nerdybill(7384)

The area of the floor in James' rectangular storage rental locker is 144 sqaure feet. The Length of the locker is 30 feet less than six times the width. What are the dimensions of the floor?
.
Let w = width of floor
then
6w-30 = length of floor
.
w(6w-30) = 144
6w^2 - 30w = 144
w^2 - 5w = 24
w^2 - 5w - 24 = 0
(w-8)(w+3) = 0
w = {-3, 8}
We can toss out the negative solution leaving only that:
w = 8 feet (width)
.
Length:
6w-30 = 6(8)-30 = 48-30 = 18 feet (length)
.
dimensions of floor: 8x18 feet